OpenStudy (anonymous):

Can Anyone Explain How To Do This Problem Please?! 17^C 4 Ill write it in the reply

OpenStudy (anonymous):



OpenStudy (anonymous):

????

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 17 C 4 = (17!)/(4!*(17-4)!) 17 C 4 = (17!)/(4!*13!) 17 C 4 = (17*16*15*14*13!)/(4!*13!) 17 C 4 = (17*16*15*14)/(4!) 17 C 4 = (17*16*15*14)/(4*3*2*1) 17 C 4 = (57120)/(24) 17 C 4 = 2380

OpenStudy (anonymous):

ooohhhh I get t thank you so much!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

do you do the same thing even if its a different letter?

jimthompson5910 (jim_thompson5910):

If it's a P instead of a C, then no

OpenStudy (anonymous):

Ok so how would you work a P Example 12^P 5

jimthompson5910 (jim_thompson5910):

n P r = (n!)/((n-r)!) 12 P 5 = (12!)/((12-5)!) 12 P 5 = (12!)/(7!) 12 P 5 = (12*11*10*9*8*7!)/(7!) 12 P 5 = 12*11*10*9*8 12 P 5 = 95040

OpenStudy (anonymous):

Ok thank you you tought it better than my teacher!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Ok so how would you work a P Example 12^P 5